2) A metal bucket of a height of 0.600 m and a mass of 450 kg is suspended from a rope in an open tank containing a liquid having a density equal to 1030 kg / m 3. The block is submerged 0100 m below the surface Calculate:

A) The force upwards in the bottom of the block.

B) the tensioning force of the rope.

C) the magnitude of the buoyant force in the cube using the Archimedes principle.
Answer #1

Given :

Height of block h = 0.6 m

Mass of block m= 450 kg

Density of liquid \rho= 1030 kg/m3

Submerged height h'= 0.10 m

A) the upward force F = pressure × surface area (bottom)

F = 1030× 9.81×0.10× 3.14× 0.22 = 127 N

B)

Tension force of the rope T = mg

T = 450×9.81 = 4414.5 N

C) according to Archimedes principle:

Magnitude of buoyant force = weight of fluid displaced by body

B = density of fluid × volume displaced

= 1030× area× 0.10